when they say find the equivalent resistan


when they say find the equivalent resistance between point A and B do they mean the total R effective, by using R1+R2+R3 for series and 1/R1+1/R2+1/R3 for parellel?
Answer #1

Yes you are right. Equivalent resistance in the given case is simply the series sum for part 1 and parallel sum for part 2

1. Req = R1 +R2 + R3 = 220 + 180 + 470 = 870 Ω

2. 1/Req = 1/R1 + 1/R2 + 1/R3  = 1/220 + 1/180 + 1/470

or Req = 220*180*470/(180*470 + 220*470 + 220*180) = 81.8 Ω